OpenStudy (anonymous):

Evaluate \[\int\limits_{C}^{}xy dy-y ^{2}dx\] where C is the closed curve consisting of the top half of the circle\[x^{2}+y ^{2}=4\] and the x-axis segment \[-2 \le x \le 2.\]

OpenStudy (anonymous):

@Zarkon Any ideas?

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

the more important question is ...do you have any ideas?

OpenStudy (anonymous):

Yes. I know \[C=C _{i}=\] \[i=1: x=2cost, y=2sint, 0 \le t \le \pi\] \[i=2: x=t, y=0, -2 \le t \le 2\] \[\int\limits_{C}^{} F \dot \times dr=\int\limits_{C _{1}}^{}F \dot \times dr _{1} +\int\limits_{C}^{}F \dot \times dr _{2}\]

OpenStudy (anonymous):

The times with the dot over it is supposed to be the dot product, I couldn't find the button for it.

OpenStudy (zarkon):

\[F\cdot dr\] F\cdot dr

OpenStudy (anonymous):

Oh thanks

OpenStudy (zarkon):

replace x and y...then compute dy and dx and replace them too...then integrate or you can use Green's theorem

OpenStudy (zarkon):

\[\int\limits_{C_1}\left[2\cos(t)2\sin(t)\frac{d(2\sin(t))}{dt}-(2\sin(t)) ^{2}\frac{d(2\cos(t))}{dt}\right]dt\]

OpenStudy (anonymous):

I've got that part, but what about the \[\cdot dr\] part of the integral?

OpenStudy (zarkon):

you don't need to do that

OpenStudy (zarkon):

your original question was not written in terms of vector fields

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

So would the correct formula be just\[\int\limits_{C}^{}fds=\int\limits_{C _{1}}^{}fds _{1}+\int\limits_{C _{2}}^{}fds _{2}\]?

OpenStudy (anonymous):

You are integrating a differential form over a manifold. So the idea is to come up with coordinate patches (parameterizations) that cover the manifold, pull the form back on those parameterizations and integrate the pulled back form (for each parameterization) and sum them together.

OpenStudy (anonymous):

Huh??

OpenStudy (zarkon):

lol

OpenStudy (zarkon):

do the integral I suggested above.

OpenStudy (zarkon):

for the upper half of the circle \[\int\limits_{C}xy dy-y ^{2}dx=\int\limits_{0}^{\pi}\left[x(t)y(t) \frac{dy}{dt}-y(t) ^{2}\frac{dx}{dt}\right]dt\]

OpenStudy (anonymous):

I guess differential geometry isn't really discussed in the class you are taking. Please ignore what I said and listen to Zarkon.